OpenStudy (anonymous):

Find all orders pairs (x, y) that satisfy the system:

OpenStudy (anonymous):

\[x ^{3}-x ^{2}y+5xy ^{2}-y ^{3}=124\] \[x ^{3}-5x ^{2}y+xy ^{2}-y ^{3}=4\]

hero (hero):

4x^2y + 4xy^2 = 120 4xy(x + y) = 120 xy(x + y) = 30 (x,y) = (3,2) (x,y) = (2,3)

hero (hero):

@CrystalSkull

OpenStudy (anonymous):

How do you get that from xy(x+y)=30?

hero (hero):

Because once you get to that point it becomes this: You know 6 x 5 = 30 Therefore make xy = 6 and (x + y) = 5 Now you're looking for two numbers that multiply to get 6 but add to get 5. You do that all the time when solving quadratic equations.

OpenStudy (anonymous):

k thanks

hero (hero):

What grade are you in?

OpenStudy (anonymous):

11th

hero (hero):

I see. Well, there could be another way to do it, but I'm pretty sure this is the easiest.

hero (hero):

I'd be interested in knowing what your teacher's solution is to this.

hero (hero):

To see if he or she does it differently

OpenStudy (anonymous):

I'll remember to post it if he ever goes over it.

hero (hero):

OK, deal. At the very least, just ask him for an alternative solution to the one you have.

hero (hero):

I'm really just interested in knowing if his is easier than mine. I doubt if it is.